Assume that your parents wanted to have $100,000 saved for college by your 18th birthday and they started saving on your first birthday. They saved the same amount each year on your birthday and earned 10.5% per year on their investments. a. How much would they have to save each year to reach their​ goal? b. If they think you will take five years instead of four to graduate and decide to have $140,000 saved just in​ case, how much would they have to save each year to reach their new​ goal? a. How much would they have to save each year to reach their​ goal? To reach the goal of $100,000​, the amount they have to save each year is
Answer #1

a) FV = 100,000
rate = 10.5%
Number of Years = 18
Amount to be saved each year = FV/((1+r)n -1)/r = 100,000/((1+10.5%)18-1)/10.5% = 2086.30

b) FV = 140,000
rate = 10.5%
Number of Years = 18
Amount to be saved each year = FV/((1+r)n -1)/r = 140,000/((1+10.5%)18-1)/10.5% = 2920.82

Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well